# Integral modelling and Reinforcement Learning control of 3D liquid metal coating on a moving substrate

### v1

@@ VIII Conclusion and Perspectives
In this study, we developed a 3D Integral Boundary Layer model to describe a liquid film on a moving substrate under the effect of a transverse magnetic field. The model was implemented in a numerical environment using the Fourier pseudo-spectral method. To reproduce open-domain conditions, we applied a linear, perfectly matched layer around the boundaries to absorb the incoming waves. Leveraging this numerical environment, we addressed a control problem involving unstable waves interacting with gas-jet and electromagnetic actuators. The optimal control law for the actuators was found using a reinforcement learning algorithm called proximal policy optimisation (PPO).
The Boundary Layer model exhibits good consistency with the non-magnetic model. As the magnetic field intensity decreases, the flat film steady-state solution and the closure relations for the flux terms and wall shear stress converge to their non-magnetic solutions.
@@
Our analysis of the unsteady model at the leading order revealed that the convective velocity is always negative for $\mbox{Ha}>2$. This highlights the strong influence that the magnetic field could have on the absolute and convective nature of steady-state solutions, further investigation of which could provide valuable insight into whether small defects on the free surface could impact industrial production processes. In addition, our study suggests that an effective magnetic field requires $\mbox{Ha}=16$ for galvanising conditions, which may be impractical for real-world applications due to high thermal gradients caused by the Joule effect and the destabilising magnetic effects on the strip’s dynamics.
Concerning the control problem, this study shows that an electromagnetic feedback regulator can effectively reduce unwanted wave amplitudes by raising their valleys. This outcome was achieved through a control mechanism governed by the resistive effects of the Lorentz force. These results open new avenues for applying this technique to more complex scenarios, particularly when considering the heat effects induced by the electromagnetic field, which could help to stabilise the liquid film via the Marangoni effects.
@@
A key feature of our model is the absence of a hysteresis term in the magnetic field. In practice, the magnetic field does not change instantaneously when $b_{t}$ is adjusted; rather, it exhibits a response time determined by the magnetic properties of the solenoid’s core material. Although this simplification is sufficient for the accuracy required in this study, a more advanced model would need to account for the hysteresis effect. Another promising direction for future research could involve examining the effects of a high-frequency magnetic field, which could generate a ‘skin effect’, concentrating the induced electric current in a small region near the free surface.
@@
This investigation provides valuable insights into the dynamics of the coating process, capturing the intricate multiscale interactions between the liquid zinc and control actuators at a computational cost lower than that of Large Eddy Simulations (LES). Beyond its immediate applications, the model can be adapted to other engineering problems. This opens new avenues for modelling, experimental investigations, and control design, ranging from ‘liquid walls’ used to confine plasma in tokamak nuclear fusion reactors to shaping versatile circuitry for next-generation electronic devices.
@@ Appendix A Magnetohydrodynamic governing equations and boundary conditions
where $\mathbf{f}_{L}$ is the Lorentz force given by the interaction between the magnetic field in the liquid and the induced current. The relative motion of the liquid zinc with respect to the external magnetic field generates an induced current in the bulk $\mathbf{j}=(j_{x},j_{y},j_{z})^{T}$, given by the Faraday’s law of induction neglecting the electric potential difference [37]:
@@
$[\bm{\tau}\mathbf{n}-\bm{\tau}_{m}\mathbf{n}]^{\rm l}_{\rm g}=2\sigma K(h),$ (63)
where $K(h)$ is the free surface mean curvature [16, Chapter 2], $[f]^{\rm l}_{\rm g}=f-f_{\rm g}$ defines the jump at the interface between the liquid and the gas phases, $\bm{\tau}$ is the viscous stress tensor, and $\bm{\tau}_{m}$ is the Maxwell stress tensor [55, Section 3.9] defined as:
@@
where $\mathbf{E}$ is the strain-rate tensor, we obtain the three scalar equations representing the projections of the force balance (63) in the local Cartesian reference frame (O;$n,\tau_{x},\tau_{z}$) (defined in (2)):
@@
$\begin{gathered}\frac{2\mu[\partial_{z}h\left(\partial_{x}h\left(\partial_{z}u+\partial_{x}w\right)-\partial_{z}v-\partial_{y}w\right)}{(\partial_{z}h)^{2}+(\partial_{x}h)^{2}+1}\\ -\frac{\partial_{x}h\left(\partial_{y}u+\partial_{x}v\right)+(\partial_{x}h)^{2}\partial_{x}u]+(\partial_{z}h)^{2}\partial_{z}w+\partial_{y}v)}{(\partial_{z}h)^{2}+(\partial_{x}h)^{2}+1}\\ +p_{g}=\frac{\sigma\left[\partial_{zz}h\left((\partial_{x}h)^{2}+1\right)-2\partial_{z}h\partial_{x}h\partial_{xz}h\right]}{\left[(\partial_{z}h)^{2}+(\partial_{x}h)^{2}+1\right]^{3/2}}\\ +\frac{\sigma\left(\left((\partial_{z}h^{2}+1\right)\partial_{xx}h\right)}{\left((\partial_{z}h)^{2}+(\partial_{x}h)^{2}+1\right)^{3/2}}+p\\ -\frac{b^{2}(\chi_{g}-\chi)\left[(\partial_{z}h)^{2}+(\partial_{x}h)^{2}-1\right]}{2\mu_{M}(\chi_{g}+1)\left[(\partial_{z}h)^{2}+(\partial_{x}h)^{2}+1\right]}\,,\end{gathered}$ (67a)
$\begin{gathered}\tau_{g,x}+\frac{\mu\left(\partial_{z}h\left(\partial_{x}h\left(\partial_{z}v+\partial_{y}w\right)+\partial_{z}u+\partial_{x}w\right)\right)}{\sqrt{(\partial_{x}h)^{2}+1}\sqrt{(\partial_{z}h)^{2}+(\partial_{x}h)^{2}+1}}\\ +\frac{\mu\left(2\partial_{x}h\left(\partial_{x}u-\partial_{y}v\right)\right)}{\sqrt{(\partial_{x}h)^{2}+1}\sqrt{(\partial_{z}h)^{2}+(\partial_{x}h)^{2}+1}}\\ +\frac{\mu\left(\left((\partial_{x}h)^{2}-1\right)\partial_{y}u+\left((\partial_{x}h)^{2}-1\right)\partial_{x}v\right)}{\sqrt{(\partial_{x}h)^{2}+1}\sqrt{(\partial_{z}h)^{2}+(\partial_{x}h)^{2}+1}}\\ +\frac{b^{2}(\chi-\chi_{g})\partial_{x}h}{\mu(\chi_{g}+1)\sqrt{(\partial_{x}h)^{2}+1}\sqrt{(\partial_{z}h)^{2}+(\partial_{x}h)^{2}+1}}=0\,,\end{gathered}$ (67b)
@@ Appendix B Approximated magnetic field of a circular finite-length solenoid
We consider a circular solenoid with radius $R_{s}$ and length $L_{s}$, as illustrated in Figure 4, with a cylindrical reference frame $\mathcal{O}(r_{s},z_{s},\varphi_{s})$ centred at its midpoint. The analytical expression for the axial $b_{za}$ and radial $b_{ra}$ components of the generated magnetic field read [43]:
@@
Figure 22: Normalized solenoid magnetic field $|\mathbf{b}|/b_{0}$ with magnetic lines (white continuous lines) in a axial-radial plane scaled with the solenoid radius $R_{s}$ and lengths $L_{s}$.
Figure 23: (a) Axia and (b) radial components of the magnetic induction field in the reduced radial direction $r_{s}/R_{s}$ at different reduced axial locations $z_{s}/L_{s}$.
Figure 22 shows the normalized magnetic field (colour map) and magnetic field lines (white continuous lines), and figure and 23 shows the radial distributions of (a) the axial $b_{za}$ and (b) radial $b_{ra}$ components of the magnetic field normalized with $b_{0}$ at different axial positions, $z_{s}/L_{s}$, outside the solenoid. The field intensity drops sharply outside the solenoid, falling below $3\%$ of $b_{0}$ for $r_{s}/R_{s}>1.5$ and $z_{s}/L_{s}>1$. Near the solenoid exit ($z_{s}/L_{s}=0.8$), $b_{za}$ remains nearly constant along $r_{s}$, while $b_{ra}$ exhibits a peak at the edge ($r_{s}/R_{s}=1$), where the field lines are most curved. As $z_{s}$ increases, both $b_{za}$ and $b_{ra}$ become smoother, adopting a bell-shaped profile in $r_{s}$, with the axial component dominating, being one order of magnitude larger than the radial component.
@@ Appendix C Governing and auxiliary equations implemented in the spectral BLEW 3D environment
$\begin{split}\partial_{\hat{t}}\hat{q}_{\hat{x}}+\zeta_{3}+\zeta_{4}=\delta^{-1}\Big{[}\hat{h}(-\zeta_{5}+\zeta_{8}+\zeta_{11}+1)\\ -H_{a}^{2}\hat{b}^{2}\hat{q_{x}}+\Delta\tau_{\hat{x}}\Big{]}\end{split}$ (80b)
$\begin{split}\partial_{\hat{t}}\hat{q}_{\hat{z}}+\zeta_{12}+\zeta_{13}=\delta^{-1}\Big{[}\hat{h}(-\zeta_{14}+\zeta_{15}+\zeta_{16})\\ -H_{a}^{2}\hat{b}^{2}\hat{q_{z}}+\Delta\tau_{\hat{z}}\Big{]}\end{split}$ (80c)
@@ Acknowledgements.
F. Pino was supported by an F.R.S.-FNRS FRIA grant. This project was funded by Arcelor-Mittal Research. B. Scheid is Research Director at F.R.S.-FNRS.

### v2

@@ VIII Conclusion and Perspectives
In this study, we developed a 3D Integral Boundary-Layer model to describe a liquid film on a moving substrate subjected to a transverse magnetic field in the context of hot-dip galvanising. The model was implemented in a numerical environment using the Fourier pseudo-spectral method. To reproduce open-domain conditions, we applied a linear, perfectly matched layer around the boundaries to absorb the incoming waves. Leveraging this numerical environment, we addressed a control problem involving unstable waves interacting with gas-jet and electromagnetic actuators. The optimal control law for the actuators was determined using the proximal policy optimisation (PPO) reinforcement learning algorithm.
The Integral Boundary Layer (IBL) model is consistent with the model without magnetic forces: as the magnetic field intensity decreases, the flat-film steady-state solution and the closure relations for the flux terms and wall shear stress converge to their well-known counterparts in the absence of magneto-hydrodynamic effects. It should be noted, however, that the investigated conditions at a reduced Reynolds number $\delta=266$, representative of galvanising flows, lie beyond the strict asymptotic range in which the model is formally derived. Yet, since previous studies have demonstrated that such models are predictive well beyond this limit, we believe that the present formulation still captures the essential coupling between hydrodynamic and electromagnetic effects. Accordingly, this work should be regarded not as an industrially deployable solution, but as a proof of concept demonstrating the feasibility of combining reduced-order modelling and reinforcement learning for the control of magnetohydrodynamic coating flows.
@@
Our analysis of the unsteady model at the leading order revealed that the convective velocity is always negative for $\mbox{Ha}>2$. This highlights the strong influence of the magnetic field on the absolute and convective nature of steady-state solutions, and further investigation could provide valuable insight into whether small defects on the free surface affect industrial production processes. In addition, our study suggests that an effective magnetic field requires $\mbox{Ha}=16$ for galvanising conditions, which may be impractical for real-world applications due to high thermal gradients caused by the Joule effect and the destabilising magnetic effects on the strip’s dynamics.
Regarding the control problem, this study demonstrates that an electromagnetic feedback regulator can effectively reduce the amplitude of unwanted waves by raising their valleys. This effect is achieved through a control mechanism driven by the resistive influence of the Lorentz force. From a practical implementation perspective, control performance against undulation perturbations can be enhanced by arranging multiple pairs in series along the streamwise direction, thereby leveraging the translational invariance of the governing equations.
@@
These findings open new possibilities for extending this approach to more complex scenarios, particularly when incorporating the thermal effects generated by the electromagnetic field, which may help stabilise the liquid film via Marangoni forces. A notable feature of our model is the absence of a hysteresis term in the magnetic field. In reality, the magnetic field does not respond instantaneously to changes in $b_{t}$; its response is determined by the magnetic properties of the solenoid’s core. While this simplification is adequate for the accuracy requirements of this study, a more refined model would need to include hysteresis effects. Another promising avenue for future research is exploring the impact of high-frequency magnetic fields, which can produce a ‘skin effect’ that confines the induced electric current to a thin layer near the free surface.
@@
This investigation provides valuable insights into the dynamics of the coating process, capturing the intricate multiscale interactions between the liquid zinc and control actuators at a computational cost lower than that of Large Eddy Simulations (LES). Beyond its immediate applications, the model can be adapted to other engineering problems. This opens new avenues for modelling, experimental investigations, and control design, ranging from ‘liquid walls’ used to confine plasma in tokamak nuclear fusion reactors to the shaping of versatile circuitry for next-generation electronic devices.
@@ Appendix A Magnetohydrodynamic governing equations and boundary conditions
where $\mathbf{f}_{L}$ is the Lorentz force given by the interaction between the magnetic field in the liquid and the induced current. The relative motion of the liquid zinc with respect to the external magnetic field generates an induced current in the bulk $\mathbf{j}=(j_{x},j_{y},j_{z})^{T}$, given by Faraday’s law of induction, neglecting the electric potential difference [14]:
@@
$\bm{\tau}\mathbf{n}-\bm{\tau}_{m}\mathbf{n}]^{\rm l}_{\rm g}=2\sigma K(h),$ (69)
where $K(h)$ is the free surface mean curvature [24, Chapter 2], $f]^{\rm l}_{\rm g}=f-f_{\rm g}$ defines the jump at the interface between the liquid and the gas phases, $\bm{\tau}$ is the viscous stress tensor, and $\bm{\tau}_{m}$ is the Maxwell stress tensor [10, Section 3.9] defined as:
@@
[…]
@@
[…]
@@ Appendix B Approximated magnetic field of a circular finite-length solenoid
We consider a circular solenoid with radius $R_{s}$ and length $L_{s}$, with a cylindrical reference frame $\mathcal{O}(r_{s},z_{s},\varphi_{s})$ centred at its midpoint. The analytical expression for the axial $b_{za}$ and radial $b_{ra}$ components of the generated magnetic field read [20]:
@@
Figure 19: Normalised solenoid magnetic field $|\mathbf{b}|/b_{0}$ with magnetic lines (white continuous lines) in an axial-radial plane scaled with the solenoid radius $R_{s}$ and lengths $L_{s}$.
Figure 20: (a) Axial
and (b) radial components of the magnetic induction field in the reduced radial direction $r_{s}/R_{s}$ at different reduced axial locations $z_{s}/L_{s}$.
Figure 19 shows the normalized magnetic field (colour map) and magnetic field lines (white continuous lines), and figure and 20(b) shows the radial distributions of (a) the axial $b_{za}$ and (b) radial $b_{ra}$ components of the magnetic field normalized with $b_{0}$ at different axial positions, $z_{s}/L_{s}$, outside the solenoid. The field intensity drops sharply outside the solenoid, falling below $3\%$ of $b_{0}$ for $r_{s}/R_{s}>1.5$ and $z_{s}/L_{s}>1$. Near the solenoid exit ($z_{s}/L_{s}=0.8$), $b_{za}$ remains nearly constant along $r_{s}$, while $b_{ra}$ exhibits a peak at the edge ($r_{s}/R_{s}=1$), where the field lines are most curved. As $z_{s}$ increases, both $b_{za}$ and $b_{ra}$ become smoother, adopting a bell-shaped profile in $r_{s}$, with the axial component dominating, being one order of magnitude larger than the radial component.
@@ Appendix C Governing and auxiliary equations implemented in the spectral BLEW 3D environment
[…]
@@ Acknowledgements.
F. Pino was supported by an F.R.S.-FNRS FRIA grant, and E. Fracchia was supported by the Short Training Program scholarship provided by the Von Karman Institute for Fluid Dynamics. This project was funded by Arcelor-Mittal Research. B. Scheid is Research Director at F.R.S.-FNRS.